# Are all models wrong?Fundamental limits in distribution-free empirical model falsification

### v1

@@ 1 Introduction
A central goal in machine learning and statistics is to learn a rule that captures properties of interest of an unknown, data-generating distribution. For instance, in a regression setting, one might seek to identify a function whose output can accurately predict the response given an observed covariate. Typically in such a setting, one does not allow the fitted function to take any arbitrary form, but rather requires it to be chosen from a (parametric or non-parametric) family of functions, a so-called model class, $\mathcal{F}$. Such restrictions to certain model classes can have many different motivations, such as encoding prior information, aiding interpretability, or allowing computationally efficient selection of a function. Indeed, any specific choice of analysis method restricts our model class in some way. This restriction may be explicit—for instance, with linear regression, we clearly limit ourselves to linear functions. But such a restriction can also be more implicit, such as through the choice of a specific neural network architecture, which is known to restrict the set of functions that can be represented, or the employed method used to eventually select a rule from the model class (DeVore et al.,, 2021; Petrova and Wojtaszczyk,, 2023).
Choosing a model class $\mathcal{F}$ involves trading off between multiple considerations. A model class $\mathcal{F}$ that is too constrained might mean that there is no good model $f\in\mathcal{F}$ (all models in the class are too simple to capture the true distribution). On the other hand, a model class $\mathcal{F}$ that is too complex may lead to challenges in selecting a good model $f\in\mathcal{F}$, if the available sample size is too small. In statistical learning theory, this is often referred to as the approximation–estimation-tradeoff (Shalev-Shwartz and Ben-David,, 2014). Moreover, an overly rich model class $\mathcal{F}$ can also lead to computational constraints, where searching within $\mathcal{F}$ is too costly, leading to an approximation–estimation–computation-tradeoff (Bottou and Bousquet,, 2007).
@@
To make this question more precise, let $R_{P}(f)$ denote the risk of a model relative to the (unknown) data distribution $P$—for example, the misclassification probability, in the context of a classification problem. Then we would like to determine whether the quantity $R_{P}(\hat{f})-\inf_{f\in\mathcal{F}}R_{P}(f)$ (sometimes referred to as the estimation error (Devroye et al.,, 2013)) is approximately zero, without relying on untestable assumptions on $P$. In order to do so, let us examine the feasibility of estimating both quantities: $R_{P}(\hat{f})$, the risk of our fitted model, and $\inf_{f\in\mathcal{F}}R_{P}(f)$, which we will refer to as the model class risk. Indeed, the difficulty of estimation is very different for these two quantities:
@@ 2.1 Setting and notation
Given a choice of loss function $\ell$, we then define the risk of a particular model $f$ as
@@ 2.2 Key questions: lower bounds and upper bounds
While the questions of obtaining lower and upper bounds appear symmetric in the above definition, in fact they are substantially different. This is because the target of inference is an infimum, $R_{P}(\mathcal{F})=\inf_{f\in\mathcal{F}}R_{P}(f)$. Therefore, a valid upper bound on $R_{P}(\mathcal{F})$ can be obtained by simply bounding $R_{P}(f)$ for any single choice of $f\in\mathcal{F}$. For example, a common strategy would be to partition the available data $\mathcal{D}_{n}$ into two independent subsets, $\mathcal{D}_{n/2}=(Z_{1},\dots,Z_{\lceil n/2\rceil})$ and $\mathcal{D}_{n/2}^{\prime}=(Z_{\lceil n/2\rceil+1},\dots,Z_{n})$, then train a model $\hat{f}\in\mathcal{F}$ using $\mathcal{D}_{n/2}$, and finally use $\mathcal{D}_{n/2}^{\prime}$ as a holdout set to provide an unbiased estimate $\hat{R}(\hat{f},\mathcal{D}_{n/2}^{\prime})$ of the risk $R_{P}(\hat{f})$ (and can then be used to construct an upper bound, via concentration arguments).
@@
In other words, any meaningful answer to this question must have $\mathbb{P}_{P}\big{\{}\hat{L}_{\alpha}(\mathcal{F},\mathcal{D}_{n})>0\}$ substantially larger than $\alpha$; we will therefore refer to any lower bound with $\mathbb{P}_{P}\big{\{}\hat{L}_{\alpha}(\mathcal{F},\mathcal{D}_{n})>0\}\lessapprox\alpha$ as trivial. Thus, the central question of this work is to determine whether it is possible to construct a lower bound on $R_{P}(\mathcal{F})$ that is always valid, and is also nontrivial, to ensure that it is more informative than the meaningless solution constructed in (2).
@@ 2.3 Overview of contributions
    The high-complexity regime. At the other extreme, consider a highly complex model class $\mathcal{F}$, which is able to interpolate not just our $n$ training points but is in fact able to interpolate a data set of size $\gg n^{2}$. Then any valid distribution-free lower bound must necessarily be trivial, with $\mathbb{P}\{\hat{L}_{\alpha}(\mathcal{F},\mathcal{D}_{n})>0\}\lessapprox\alpha$.
@@
    The in-between regime. Finally, we find an interesting regime in between these two extremes. If $\mathcal{F}$ interpolates a data set of size $n$, but its interpolation capacity is $\lesssim n^{2}$, then the question does not have a general answer: we will see specific examples of model classes with qualitatively different behavior within this regime. Counterintuitively, it is sometimes possible to verify that $R_{P}(\mathcal{F})>0$ (“all models are wrong”) even when $\hat{R}(\mathcal{F},\mathcal{D}_{n})=0$ (i.e., the empirical risk does not provide any evidence that “all models are wrong”).
@@ 3.1 Defining interpolation capacity
As summarized in Figure 1, the low-complexity and high-complexity regimes are characterized by the interpolation capacity of the model class $\mathcal{F}$—informally, how large of a sample can be interpolated by some $f\in\mathcal{F}$. Before presenting our theoretical results, we begin by formalizing this measure of the complexity of $\mathcal{F}$. These definitions are closely related to complexity measures such as VC dimension and fat-shattering dimension that appear in the statistical learning theory literature (Devroye et al.,, 2013).
@@
By definition, we have $N(\mathcal{F},P)\leq N_{+}(\mathcal{F},P)$—and in fact, in many common examples, these two measures of interpolation capacity are equal or approximately equal. For instance, in a regression setting with data points $(X_{i},Y_{i})\in\mathbb{R}^{d}\times\mathbb{R}=:\mathcal{Z}$, if $\mathcal{F}$ is the class of all linear models (without an intercept), then $N(\mathcal{F},P)=N_{+}(\mathcal{F},P)=d$ for any distribution $P$ with a density on $\mathbb{R}^{d}\times\mathbb{R}$. For this reason, when discussing our results, we treat these two different measures of interpolation capacity as essentially interchangeable, for the purpose of intuition.
3.2 A lower bound via empirical risk for the low-complexity regime
@@
If $R_{P}(\mathcal{F})=0$ then we must have $\hat{R}(\mathcal{F},\mathcal{D}_{n})=0$ almost surely. If instead $R_{P}(\mathcal{F})>0$, then by Markov’s inequality, we have $\mathbb{P}_{P}\{\hat{L}_{\alpha}^{\mathrm{ERM}}(\mathcal{F},\mathcal{D}_{n})\leq R_{P}(\mathcal{F})\}=\mathbb{P}_{P}\{\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq\alpha^{-1}R_{P}(\mathcal{F})\}\geq 1-\alpha$, since $\mathbb{E}_{P}[\hat{R}(\mathcal{F},\mathcal{D}_{n})]\leq R_{P}(\mathcal{F})$ as computed in (3). ∎
@@ 3.3 Hardness results in the high-complexity regime
To summarize what we have seen for the low-complexity and high-complexity regimes in terms of the interpolation capacity, if $\mathcal{F}$ can interpolate data sets of size $N\gg n^{2}$, then $\mathbb{P}_{P}\{\hat{L}_{\alpha}(\mathcal{F},\mathcal{D}_{n})>0\}\lessapprox\alpha$ for every valid lower bound $\hat{L}_{\alpha}(\mathcal{F},\cdot)$. Recalling the trivial solution in (2), this means that there do not exist any nontrivial valid lower bounds. On the other hand, if $\mathcal{F}$ cannot interpolate data sets of size $n$, then there always exists a positive valid lower bound, namely, $\hat{L}_{\alpha}^{\mathrm{ERM}}(\mathcal{F},\cdot)$.
@@ 4 Examples at the boundaries: the in-between regime
  • For one example (piecewise constant functions), we will see that a nontrivial lower bound is possible whenever $N(\mathcal{F},P)\lesssim n^{2}$—and in particular, we may have $N(\mathcal{F},P)\geq n$.
@@ 4.2 Example: linear models
We first need to set up some notation and definitions. To begin, let us define a class of noise-free linear distributions,
@@
Let $\tilde{\mathcal{Q}}^{(n)}_{\textnormal{lin}}$ denote the family of mixtures of distributions in $\mathcal{Q}^{(n)}_{\textnormal{lin}}$, which can be seen as the convex hull of $\mathcal{Q}^{(n)}_{\textnormal{lin}}$. Finally, define
@@
We can therefore see that any valid lower bound is essentially trivial whenever $\lambda_{n,d}(P)\approx 0$—but when will this be the case? The next result establishes that, in a high-dimensional setting where $d\geq n$, the quantity $\lambda_{n,d}(P)$ is small for a broad class of distributions $P$.
@@
Let $P$ be a distribution on $\mathbb{R}^{d}\times\mathbb{R}$ with a density. Let $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P$, and define $\mathbf{X}\in\mathbb{R}^{n\times d}$ as the matrix with rows $X_{i}$. If $d\geq n$, then for any positive definite $\Omega\in\mathbb{R}^{d\times d}$,
@@
Proposition 8 suggests that as long as $h_{\Omega}(\mathbf{X})=(\det(\mathbf{X}\Omega\mathbf{X}^{\top}))^{-1/2}$ concentrates for some positive definite $\Omega$, we expect $\lambda_{n,d}(P)\approx 0$. The next result derives an explicit upper bound on $\lambda_{n,d}(P)$ for the special case of a Gaussian distribution. [^4] We show in the Appendix how this type of bound can be extended to more general distributions.
@@
In particular, we have $\lambda_{n,d}(P)\approx 0$ for any $d\gg n$ in the Gaussian setting, meaning that any valid lower bound on the model class risk of $\mathcal{F}_{\mathrm{lin}}^{(d)}$ is essentially trivial. This result therefore offers another example for the “in-between” regime of Figure 1, which complements the findings of Section 4.1 for our first example: it verifies that the low-complexity regime results of Theorem 2 and Corollary 3 are essentially tight, in the sense that if $N(\mathcal{F},P)\gg n$, then we cannot provide a universal guarantee that a nontrivial lower bound is always possible for any $\mathcal{F}$.

### v2

@@ 1 Introduction
A central goal in machine learning and statistics is to learn a rule that captures properties of interest of an unknown, data-generating distribution. For instance, in a regression setting, we might seek to identify a function whose output can accurately predict the response given an observed covariate. Typically in such a setting, we do not allow the fitted function to take any arbitrary form, but rather require it to be chosen from a (parametric or non-parametric) family of functions, a so-called model class, $\mathcal{F}$. Such restrictions to certain model classes can have many different motivations, such as encoding prior information, aiding interpretability, or allowing computationally efficient selection of a function. Indeed, any specific choice of analysis method restricts our model class in some way. This restriction may be explicit—for instance, with linear regression, we clearly limit ourselves to linear functions. But such a restriction can also be more implicit, such as through the choice of a specific neural network architecture, which is known to restrict the set of functions that can be represented, or the employed method used to eventually select a rule from the model class (DeVore et al.,, 2021; Petrova and Wojtaszczyk,, 2023).
Choosing a model class $\mathcal{F}$ involves trading off between multiple considerations. A model class $\mathcal{F}$ that is too constrained might mean that there is no good model $f\in\mathcal{F}$ (all models in the class are too simple to capture the true distribution). On the other hand, a model class $\mathcal{F}$ that is too complex may lead to challenges in selecting a good model $f\in\mathcal{F}$, if the available sample size is too small. In statistical learning theory, this is often referred to as the approximation–estimation tradeoff (Shalev-Shwartz and Ben-David,, 2014). Moreover, an overly rich model class $\mathcal{F}$ can also lead to computational constraints, where searching within $\mathcal{F}$ is too costly, leading to an approximation–estimation–computation tradeoff (Bottou and Bousquet,, 2007).
@@
To make this question more precise, let $R_{P}(f)$ denote the risk of a model relative to the (unknown) data distribution $P$—for example, the misclassification probability, in the context of a classification problem. Then we would like to determine whether the quantity $R_{P}(\hat{f})-\inf_{f\in\mathcal{F}}R_{P}(f)$ (sometimes referred to as the estimation error (Devroye et al.,, 1996)) is approximately zero, without relying on untestable assumptions on $P$. In order to do so, let us examine the feasibility of estimating both quantities: $R_{P}(\hat{f})$, the risk of our fitted model, and $\inf_{f\in\mathcal{F}}R_{P}(f)$, which we will refer to as the model class risk. Indeed, the difficulty of estimation is very different for these two quantities:
@@ 2.1 Setting and notation
Given a choice of loss function $\ell$, we define the risk of a particular model $f$ as
@@ 2.2 Key questions: lower bounds and upper bounds
While the questions of obtaining lower and upper bounds appear symmetric in the above definition, in fact they are substantially different. This is because the target of inference is an infimum, $R_{P}(\mathcal{F})=\inf_{f\in\mathcal{F}}R_{P}(f)$. Therefore, a valid upper bound on $R_{P}(\mathcal{F})$ can be obtained by simply bounding $R_{P}(f)$ for any single choice of $f\in\mathcal{F}$. For example, a common strategy would be to partition the available data $\mathcal{D}_{n}$ into two independent subsets, $\mathcal{D}_{n/2}=(Z_{1},\dots,Z_{\lceil n/2\rceil})$ and $\mathcal{D}_{n/2}^{\prime}=(Z_{\lceil n/2\rceil+1},\dots,Z_{n})$, then train a model $\hat{f}\in\mathcal{F}$ using $\mathcal{D}_{n/2}$, and finally use $\mathcal{D}_{n/2}^{\prime}$ as a holdout set to provide an unbiased estimate $\hat{R}(\hat{f},\mathcal{D}_{n/2}^{\prime})$ of the risk $R_{P}(\hat{f})$ (and we can also construct an upper bound on $R_{P}(\hat{f})$, via concentration arguments).
@@
In other words, any meaningful answer to this question must have $\mathbb{P}_{P}\big{\{}\hat{L}_{\alpha}(\mathcal{F},\mathcal{D}_{n})>0\}$ substantially larger than $\alpha$; we will therefore refer to any lower bound with $\mathbb{P}_{P}\big{\{}\hat{L}_{\alpha}(\mathcal{F},\mathcal{D}_{n})>0\}\leq\alpha+\mathrm{o}(1)$ as trivial. Thus, the central question of this work is to determine whether it is possible to construct a lower bound on $R_{P}(\mathcal{F})$ that is always valid, and is also nontrivial, to ensure that it is more informative than the meaningless solution constructed in (2).
@@ 2.3 Overview of contributions
    The high-complexity regime. At the other extreme, consider a highly complex model class $\mathcal{F}$, which is able to interpolate not just our $n$ training points but is in fact able to interpolate a data set of [^3] For $a,b\geq 0$, we write $a\gg b$ in informal descriptions of our results whenever we mean that $b/a=\mathrm{o}(1)$. size $\gg n^{2}$. Then any valid distribution-free lower bound must necessarily be trivial, with $\mathbb{P}\{\hat{L}_{\alpha}(\mathcal{F},\mathcal{D}_{n})>0\}\leq\alpha+\mathrm{o}(1)$.
@@
    The in-between regime. Finally, we find an interesting regime in between these two extremes. If $\mathcal{F}$ interpolates a data set of size $n$, but its interpolation capacity is $\mathcal{O}(n^{2})$, then the question does not have a general answer: we will see specific examples of model classes with qualitatively different behavior within this regime. Counterintuitively, it is sometimes possible to verify that $R_{P}(\mathcal{F})>0$ (“all models are wrong”) even when $\hat{R}(\mathcal{F},\mathcal{D}_{n})=0$ (i.e., the empirical risk does not provide any evidence that “all models are wrong”).
@@ 3.1 Defining interpolation capacity
As summarized in Figure 1, the low-complexity and high-complexity regimes are characterized by the interpolation capacity of the model class $\mathcal{F}$—informally, how large of a sample can be interpolated by some $f\in\mathcal{F}$. Before presenting our theoretical results, we begin by formalizing this measure of the complexity of $\mathcal{F}$. These definitions are closely related to complexity measures such as VC dimension and fat-shattering dimension that appear in the statistical learning theory literature (Devroye et al.,, 1996).
@@
By definition, we have $N(\mathcal{F},P)\leq N_{+}(\mathcal{F},P)$—and in fact, in many common examples, these two measures of interpolation capacity are equal or approximately equal. For instance, in a regression setting with data points $(X_{i},Y_{i})\in\mathbb{R}^{d}\times\mathbb{R}=:\mathcal{Z}$, if $\mathcal{F}$ is the class of all linear models (without an intercept), then $N(\mathcal{F},P)=N_{+}(\mathcal{F},P)=d$ for any distribution $P$ with a density on $\mathbb{R}^{d}\times\mathbb{R}$. For this reason, when discussing our results, we treat these two different measures of interpolation capacity as essentially interchangeable, for the purpose of intuition. [^4] There are however also cases in which $N(\mathcal{F},P)\ll N_{+}(\mathcal{F},P)$—for instance, consider the classification setting from Section 2.1 with $\mathcal{F}$ containing all constant functions on $\mathbb{R}^{d}$ and $P$ being a distribution of $(X,Y)$ on $\mathbb{R}^{d}\times\{0,1\}$ such that $\mathbb{P}_{P}(Y=1|X)$ is in $(0,1)$ almost surely. Then, $N(\mathcal{F},P)=1$, while $N_{+}(\mathcal{F},P)=\infty$.
3.2 A lower bound via the empirical risk for the low-complexity regime
@@
If $R_{P}(\mathcal{F})=0$, then we must have $\hat{R}(\mathcal{F},\mathcal{D}_{n})=0$ almost surely. If instead $R_{P}(\mathcal{F})>0$, then by Markov’s inequality, we have $\mathbb{P}_{P}\{\hat{L}_{\alpha}^{\mathrm{ERM}}(\mathcal{F},\mathcal{D}_{n})\leq R_{P}(\mathcal{F})\}=\mathbb{P}_{P}\{\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq\alpha^{-1}R_{P}(\mathcal{F})\}\geq 1-\alpha$, since $\mathbb{E}_{P}[\hat{R}(\mathcal{F},\mathcal{D}_{n})]\leq R_{P}(\mathcal{F})$ as computed in (3). ∎
@@ 3.3 Hardness results in the high-complexity regime
The proof of this result follows the “sample–resample” strategy used for establishing certain hardness results in the distribution-free inference literature (see Angelopoulos et al., (2024)), which relies on the fact that, when sampling $n$ times from a population of size $N$, the total variation distance between sampling with replacement and sampling without replacement is bounded by $\frac{n^{2}}{2N}$.
@@
To summarize what we have seen for the low-complexity and high-complexity regimes in terms of the interpolation capacity, if $\mathcal{F}$ can interpolate data sets of size $N\gg n^{2}$, then $\mathbb{P}_{P}\{\hat{L}_{\alpha}(\mathcal{F},\mathcal{D}_{n})>0\}\leq\alpha+\mathrm{o}(1)$ for every valid lower bound $\hat{L}_{\alpha}(\mathcal{F},\cdot)$. Recalling the trivial solution in (2), this means that there do not exist any nontrivial valid lower bounds. On the other hand, if $\mathcal{F}$ cannot interpolate data sets of size $n$, then there always exists a positive valid lower bound, namely, $\hat{L}_{\alpha}^{\mathrm{ERM}}(\mathcal{F},\cdot)$.
@@ 4 Examples at the boundaries: the in-between regime
  • For one example (piecewise constant functions), we will see that a nontrivial lower bound is possible whenever $N(\mathcal{F},P)=\mathcal{O}(n^{2})$—and in particular, we may have $N(\mathcal{F},P)\gg n$.
@@ 4.1 Example: piecewise constant regression
The intuition for proving the above theorem is the following: consider any $f\in\mathcal{F}_{\mathrm{pwc}}^{(m)}$, which takes (at most) $m$ distinct values in its output. If $m$ is not too large, then with high probability, at most $n-1$ of these values will actually be observed in a random sample of size $n$—that is, at least one observed value will be repeated. Consequently, on the sample $\mathcal{D}_{n}$, the empirical risk of $f$ can also be achieved by some function in $\mathcal{F}_{\mathrm{pwc}}^{(n-1)}$, which is a much less complex model class.
@@ 4.2 Example: linear models
We first need to set up some notation and definitions. To begin, let us define a class of noise-free linear distributions on $\mathbb{R}^{d}\times\mathbb{R}$,
@@
Let $\tilde{\mathcal{Q}}^{(n)}_{\textnormal{lin}}$ denote the family of mixtures of distributions in $\mathcal{Q}^{(n)}_{\textnormal{lin}}$, which can be seen as the convex hull of $\mathcal{Q}^{(n)}_{\textnormal{lin}}$. Finally, for any distribution $P$ on $\mathbb{R}^{d}\times\mathbb{R}$, define
@@
For intuition, this holds simply due to the definition of total variation distance: if $\lambda_{n,d}(P)$ is low, then the distribution $P^{n}$ of the observed data is nearly indistinguishable from a convex combination of noise-free linear distributions—and consequently we cannot construct a nontrivial lower bound.
From this result, we can therefore see that any valid lower bound is essentially trivial whenever $\lambda_{n,d}(P)\approx 0$—but when will this be the case? The next result establishes that, in a high-dimensional setting where $d\geq n$, the quantity $\lambda_{n,d}(P)$ is small for a broad class of distributions $P$.
@@
Let $P$ be a distribution on $\mathbb{R}^{d}\times\mathbb{R}$ with a density. Let $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P$, and define $\mathbf{X}\in\mathbb{R}^{n\times d}$ as the matrix with rows $X_{i}$. If $d\geq n$, then for any positive definite matrix $\Omega\in\mathbb{R}^{d\times d}$,
@@
Proposition 8 suggests that as long as $h_{\Omega}(\mathbf{X})=(\det(\mathbf{X}\Omega\mathbf{X}^{\top}))^{-1/2}$ concentrates for some positive definite matrix $\Omega$, we expect $\lambda_{n,d}(P)\approx 0$. To make this intuition more precise, consider the case $\Omega=\mathbf{I}_{d}$, and observe that we can write the determinant as
$h_{\mathbf{I}_{d}}(\mathbf{X})^{-1}=\det(\mathbf{X}\mathbf{X}^{\top})^{1/2}=\|X_{1}\|_{2}\cdot\|\mathbf{P}_{X_{1}}^{\perp}X_{2}\|_{2}\cdot\|\mathbf{P}_{X_{1},X_{2}}^{\perp}X_{3}\|_{2}\cdot\ldots\cdot\|\mathbf{P}_{X_{1},\dots,X_{n-1}}^{\perp}X_{n}\|_{2},$
where for each $i$, the notation $\mathbf{P}_{X_{1},\dots,X_{i}}^{\perp}$ denotes projection to the orthogonal complement of the span of $X_{1},\dots,X_{i}\in\mathbb{R}^{d}$. Thus, as long as we expect concentration of the norms of the $X_{i}$’s, and also expect that the $X_{i}$’s are not likely to be strongly collinear, we can expect $h_{\mathbf{I}_{d}}(\mathbf{X})$ to have low variance (and so $\lambda_{n,d}(P)$ will be small).
To make this more concrete, the next result derives an explicit upper bound on $\lambda_{n,d}(P)$ for the special case of a Gaussian distribution. [^6] We show in the Appendix how this type of bound can be extended to more general distributions.
@@
In particular, we have $\lambda_{n,d}(P)\approx 0$ for any $d\gg n$ in the Gaussian setting, meaning that any valid lower bound on the model class risk of $\mathcal{F}_{\mathrm{lin}}^{(d)}$ is essentially trivial—in particular, this holds for, say, $d\propto n^{1+a}$ for any $a>0$, meaning that we may still have $d\ll n^{2}$.
This result therefore offers another example for the “in-between” regime of Figure 1, which complements the findings of Section 4.1 for our first example: it verifies that the low-complexity regime results of Theorem 2 and Corollary 3 are essentially tight, in the sense that if $N(\mathcal{F},P)\gg n$, then we cannot provide a universal guarantee that a nontrivial lower bound is always possible for any $\mathcal{F}$. A related question is also addressed by Kong and Valiant, (2019, Theorem 3), who show that even when restricted to the case where $P$ is known to be Gaussian but nothing is to be assumed about the covariance matrix, it is impossible to distinguish between $R_{P}(\mathcal{F}_{\mathrm{lin}}^{(d)})>0$ and $R_{P}(\mathcal{F}_{\mathrm{lin}}^{(d)})=0$ when given access to only $n=\mathcal{O}(d)$ data points.